OpenStudy (anonymous):

Choose the ratio that is equivalent to this one. 1/6 A. 3/8 B. 4/24 C. 2/6 D. 12/2

OpenStudy (john_es):

You can try to multiply denominator and numerator by a number (amplificate the ratio). For example, \[\frac{1}{6}\times\frac{2}{2}=\frac{2}{12}\]\[\frac{1}{6}\times\frac{3}{3}=\frac{3}{18}\] And so on.